# Resonant excitation of eccentricity in spherical extreme-mass-ratio inspirals

### v1

@@ I Introduction
EMRIs are expected to perform $10^{4}-10^{5}$ revolutions in the strong-field regime, due to their adiabatic evolution, thus will, in principle, trace the background spacetime of the primary and open a new domain of GW phenomenology [11, 12, 13] regarding the validity of general relativity (GR) [14, 15, 16, 17, 18, 19, 20], the existence of dark compact objects [21, 22, 23, 24, 25], the effect of dark matter [26, 27, 28, 29, 30, 31] and astrophysical environments around BHs [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], accretion physics [43, 44, 45], superradiant clouds [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57], ultralight fundamental fields [58, 59, 60, 61, 62, 63, 64, 65], transient orbital resonances [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85] and other relativistic effects [86].
@@
Gravitational self-force (or radiation reaction) [87, 88, 89, 90, 91, 92, 93] is one of the most successful techniques to evolve EMRIs in order to generate accurate gravitational waveforms [94, 95, 96, 97, 98, 99, 100, 101, 102]. Even so, calculating the self-force of a secondary around a Kerr supermassive primary is an extremely daunting, theory-dependent, and time-consuming task, therefore alternative methodologies have been used to approximate the orbital trajectory and resulting waveform of EMRIs. To that end, a variety of approximations to the process of radiation reaction have been constructed and used over the years, with great success [103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119].
@@
The influence of gravitational radiation reaction on orbiting particles around BHs is of great interest, especially in EMRIs [120, 121, 122, 123, 124, 125, 126, 127]. It has been shown that at the adiabatic limit, radiation reaction drives a particle in spherical motion around a Schwarzschild [120] or a Kerr BH [128, 129] through successively-damped spherical geodesics, i.e. spherical orbits of particles remain spherical (with zero eccentricity) under the action of gravitation radiation reaction up to some minimum radius which is quite close to the separatrix (the analogue of ISCO for non-equatorial orbits). This occurs due to the evolution of the Carter constant [130] and the fact that these orbits cannot satisfy the resonance condition $T_{\theta}=2nT_{r}$, for some integer $n$, where $T_{\theta}$ is the $\theta$-motion period and $T_{r}$ is the averaged period of the small-oscillation radial period [129, 128]. In spite of that, the circularity preservation of EMRIs under radiation reaction has been proven only for vacuum solutions of GR, such as the Kerr family [131, 132], which have integrable geodesics [133].
@@ IV Initial conditions and inspiral fluxes
In what follows, we will evolve integrable and non-integrable EMRIs with the technique discussed above that has been tested thoroughly in [143, 66, 67] and adapted in [74, 75, 85, 84] to include non-linear, updated fluxes every $N$ revolutions around the compact primary. Hereinafter, $N$ will vary from $10^{2}$ to $10^{3}$ in order to achieve smooth data for the evolution of the orbital elements and eventually the inspiral (see Ref. [75] for a detailed presentation on the inspiral evolution scheme).
@@ Acknowledgements.
The authors acknowledge financial support provided by the DAAD program for the “Promotion of the exchange and scientific cooperation between Greece and Germany IKYDAAD 2022” (57628320). K.D. acknowledges financial support provided under the European Union’s H2020 ERC, Starting Grant agreement no. DarkGRA–757480 and the MIUR PRIN and FARE programmes (GW-NEXT, CUP:B84I20000100001). K.D. also acknowledges hospitality from the National and Kapodistrian University of Athens, and University of Tübingen where part of this work was conducted.
@@ Appendix A Constants of motion
For the integrable system $a_{Q}=0$ the constants of motion $\{E,L_{z},\tilde{Q}\}$ can be expressed in terms of the orbital parameters: the semi-latus rectum $p$, the eccentricity $e$ and the minimum polar angle $\theta_{min}$ reached by the orbit.

### v2

@@ I Introduction
EMRIs are expected to perform $10^{4}-10^{5}$ revolutions in the strong-field regime, due to their adiabatic evolution, thus will, in principle, trace the background spacetime of the primary and open a new domain of GW phenomenology [11, 12, 13] regarding the validity of general relativity (GR) [14, 15, 16, 17, 18, 19, 20], the existence of dark compact objects [21, 22, 23, 24, 25], the effect of dark matter [26, 27, 28, 29, 30, 31] and astrophysical environments around BHs [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], accretion physics [43, 44, 45], superradiant clouds [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57], ultralight fundamental fields [58, 59, 60, 61, 62, 63, 64, 65], transient orbital and tidal resonances [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95] and other relativistic effects [96].
@@
Gravitational self-force (or radiation reaction) [97, 98, 99, 100, 101, 102, 103, 104, 105, 95] is one of the most successful techniques to evolve EMRIs in order to generate accurate gravitational waveforms [106, 107, 108, 109, 110, 111, 112, 113, 114]. Even so, calculating the self-force of a secondary around a Kerr supermassive primary is an extremely daunting, theory-dependent, and time-consuming task, therefore alternative methodologies have been used to approximate the orbital trajectory and resulting waveform of EMRIs. To that end, a variety of approximations to the process of radiation reaction have been constructed and used over the years, with great success [115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131].
@@
The influence of gravitational radiation reaction on orbiting particles around BHs is of great interest, especially in EMRIs [132, 133, 134, 135, 136, 137, 138, 139]. It has been shown that at the adiabatic limit, radiation reaction drives a particle in spherical motion around a Schwarzschild [132] or a Kerr BH [140, 141] through successively-damped spherical geodesics, i.e. spherical orbits of particles remain spherical (with zero eccentricity) under the action of gravitation radiation reaction up to some minimum radius which is quite close to the separatrix (the analogue of ISCO for non-equatorial orbits). The stability of spherical Kerr orbits has been proven by Kennefick and Ori [141] who calculated the relation between the rates of change of energy, of the axial component of angular momentum and that of the Carter constant $Q$ [142] under a generic gravitational self-force (GSF). They demonstrate that the rate of change of $Q$ has the “appropriate” form so that spherical orbits remain spherical at adiabatic order so long as the self-force does not resonate with the radial oscillations. Because the periodicity of the GSF for a spherical orbit is determined by the polar motion and due to the reflection symmetry of the metric, the frequency of GSF is twice the polar frequency. Thus the only assumption for this stability to hold is the fact that there is no resonance of the form $T_{\theta}=2nT_{r}$, for some integer $n$, where $T_{\theta}$ is the $\theta$-motion period and $T_{r}$ is the period of the small-oscillation radial period [141, 140]. This resonance condition is never met in Kerr geodesics. In spite of that, the circularity preservation of EMRIs under radiation reaction has been proven only for vacuum solutions of GR, such as the Kerr family [143, 144], which have integrable geodesics [145]. Nevertheless, the satisfaction of the resonance condition maybe occur in other metrics.
@@ IV Initial conditions and inspiral fluxes
In what follows, we will evolve integrable and non-integrable EMRIs with the technique discussed above that has been tested thoroughly in [155, 66, 67] and adapted in [83, 84, 94, 93] to include non-linear, updated fluxes every $N$ revolutions around the compact primary. More precisely, we numerically integrate the 2nd order coupled ordinary-differential equations (coupled geodesics) for a small interval of time compared to the radiation reaction timescale, to obtain $r(t)$ and $\theta(t)$, assuming a linear evolution of $E(t)$ and $L_{z}(t)$. Then we run a geodesic evolution with initial conditions taken from the final point in phase space of the previous adiabatically evolved orbit. From this geodesic orbit we determined numerically the orbital elements $(p,e,\iota)$ as
$e=\frac{r_{1}-r_{2}}{r_{1}+r_{2}},\,\,\,\,\,\,\,p=\frac{2r_{1}r_{2}}{r_{1}+r_{2}},\,\,\,\,\,\,\,\iota=\frac{\theta_{1}-\theta_{2}}{2},$ (51)
are the maximum and minimum angular values along the geodesic orbit. Finally, we use these orbital elements to update the average losses of $E$ and $L_{z}$ and we evolve again the orbit under radiation reaction for the next time interval. The time interval of each segment of adiabatic evolution has been adjusted to include $N=10^{2}-10^{3}$ orbital revolutions in order to achieve smooth data for the evolution of the orbital elements and eventually the inspiral (see Ref. [84] for a detailed presentation on the inspiral evolution scheme).
@@ V Eccentricity excitation in integrable EMRIs
We point out that all our numerical integrations, regarding the adiabatic evolution of approximately spherical orbits, have been progressed up to times where the separatrix of the specific deformed-Kerr metric has not been crossed yet. Therefore, the observed increment in eccentricity is not an artifact that could be attributed to plunging behavior of the orbit, especially because it is too sharp. Furthermore, our numerical integration scheme remains valid all the way. In contrast, the orbital characteristics at the end-point of our integrations do lie in the area beyond the separatrix of Kerr, that is in the area describing plunging orbits of Kerr. Also we were careful in order to ensure that after the resonance crossing the eccentricity was ”stabilized” for a period of time until it begins to increase again due to transition to plunge and we avoided orbits that after resonance crossing starts plunging almost immediately.
@@ Acknowledgements.
The authors acknowledge financial support provided by the DAAD program for the “Promotion of the exchange and scientific cooperation between Greece and Germany IKYDAAD 2022” (57628320). This project has received funding from the European Union’s Horizon-MSCA-2022 research and innovation programme “EinsteinWaves” under grant agreement No. 101131233. K.D. acknowledges financial support provided under the European Union’s H2020 ERC, Starting Grant agreement no. DarkGRA–757480 and the MIUR PRIN and FARE programmes (GW-NEXT, CUP:B84I20000100001). K.D. also acknowledges hospitality from the National and Kapodistrian University of Athens, and University of Tübingen where part of this work was conducted.
@@ Appendix A Constants of motion
For the integrable system $a_{Q}=0$ the constants of motion $\{E,L_{z},\tilde{Q}\}$ can be expressed in terms of the orbital parameters: the semi-latus rectum $p$, the eccentricity $e$ and the minimum polar angle $\theta_{min}$ reached by the orbit. We derive the corresponding expressions for the constants of motion in Johannsen spacetime following the method outlined in Appendix B of [152] for the Kerr metric.